OpenStudy (babyd):

i cant solve this problem 2/8/9=18/w

OpenStudy (saifoo.khan):

is that a mixed fraction?

OpenStudy (babyd):

yes

OpenStudy (anonymous):

w=8

OpenStudy (anonymous):

2/8/9 = 2*9/8 = 18/8

OpenStudy (saifoo.khan):

wanna join me in my class?

OpenStudy (babyd):

yea how do u solve this

OpenStudy (babyd):

545/8% need to convert percent into a decimal

OpenStudy (babyd):

sorry its 54 and 5/8%